# A Novel Energy-Efficient Cross-Layer Design for Scheduling and Routing in 6TiSCH Networks

### v1

@@ II-A2 Scheduling Function
Figure 5: An Example of 6P Transactions.
II-B Related Works
The conventional mechanisms, while foundational, have been scrutinized for potential limitations and areas of enhancement. This section investigates the existing literature for proposed methodologies, achievements, and gaps, thereby contextualizing the motivation and novelty for our work. Aiming to enhnace fundamental aspects of 6TiSCH, in particular the TSCH configuration, authors of [14] adopted heterogeneous time slot length in MSF to minimize communication latency. In [15], the authors introduced a heuristic for parent selection in multi-PHY TSCH, showing that slot bonding can adapt to different PHY data rates and achieve high Packet Delivery Ratios (PDRs). From the perspective of network formation in 6TiSCH, the study conducted in [16] evaluated the network formation procedure, and pinpointed numerous limitations within its scope. Then, they proffered several recommendations and established guidelines for parameter adjustments, tailoring them to several network topologies. Also, authors of [17] addressed performance degradation during 6TiSCH network formation. They highlighted the challenges faced by new nodes joining the network, primarily due to the policy of allocating only one shared cell per slotframe for control packet transmission. They particularly emphasized the limitations of prioritizing EBs during network formation. Then, they proposed an improved approach that effectively integrates control packets into the network formation process. Typically, research works that enhance SF fall into three categories: 1) Centralized approaches such as [18, 19], 2) distributed approaches like [20], and 3) autonomous approaches, e.g., [21]. In [18], the authors proposed a fully centralized schedule for data gathering, where each node has to send the allocation request to the DAG root and receive the answer, thus wasting a considerable amount of energy and time. Alternatively, in [19], a central entity is responsible for managing the scheduling of transmissions in the network and uses a recurrent neural network (RNN) based model to predict the next transmission slot for each node, aiming to reduce the latency of end-to-end communications. However, a higher control packets overhead is generated. Moreover, authors of [22] introduced an Adaptive Static Scheduling (ASS) technique designed to enhance the network’s energy efficiency. Several works delved into the scheduling strategies and network formation solutions. They are comprehensively summarized in surveys, e.g., [23, 24] Other works focused on optimizing the RPL’s objective function. For instance, authors of [25] introduced a novel parent selection function that accounts for the number of neighbors and traffic load and aiming to optimize the PDR, while minimizing frequent parent switching. In [26], a Q-learning based algorithm is proposed to refine the RPL’s objective function, while [27] proposed an early parent switching scheme, based on the remaining queue capacity of the parent node. Finally, the study in [27] introduced a novel parent switching scheme predicated on the remaining queue capacity of the parent node, which effectively improved the PDR performance.
III Proposed Solution
@@
Our work extensively focus on the intricate interplay between RPL and SF, and between SF and packet queue management. The core philosophy of our solution is as follows: 1) Reduce the transmission frequency of control packets upon parent switching; 2) Minimize the number of failed parent switching events that result from cell unavailability; 3) Prevent packet dropping events due to queue saturation; and 4) Improve the jitter and latency performances by slot selection. Our approach targets saving valuable energy resources and mitigating unnecessary medium contention, thus enhancing the spatial reuse. For understanding purposes of our solution, we highlight the following technical terms usage:
@@
  • “DAO-ACK” and “ACK” are two distinct message. DAO-ACK denotes a packet transmitted by the DAG root during a to acknowledge DAO messages for reliable route registration/de-registration. In contrast, a packet “ACK”, known also as a link-layer or TSCH acknowledgment (shown in Fig. 6), is issued upon packet reception [28].
@@
III-A Regulated Triggering of Parent Switching
In 6TiSCH networks, parent switching might be challenging. Indeed, the decision to switch to a potentially superior parent is complex, given the node’s lack of awareness regarding the new parent’s cell availability. This uncertainty can lead to inefficient energy usage since the node may continue control packet transmissions without realizing effective cell allocation. Also, the RPL specifications [29, 8] did not detail the protocol’s behavior regarding transmission interruption to the former parent during this process. Consequently, increased latency and potential packet loss may occur. To tackle this issue, our solution proposes to modify the DIO packet in order to carry a list of available slots of the sender. This approach ensures that a node initiates parent switching only when a shared available slot exists, hence improving the efficiency of the process. Algo. 1 below delineates this mechanism, particularly focusing on the node’s decision-making process at the juncture of parent switching.
@@ III-A Regulated Triggering of Parent Switching
// ------ Initialization
@@
1 $selected\_parent\leftarrow parent\mid parent.rank=\min(p.rank\textbf{ For }p\textbf{ in }parents\textbf{ and }p.free\_slot\cap child\_slots\neq\emptyset)$
2 if $selected\_parent\neq None$ then
@@ 4             $nb\leftarrow$ min_nb
6             $nb\leftarrow$ nb_parentSlot if $nb>max\_nb$ then
@@ 7                   $nb\leftarrow max\_nb$
8
9      $free\_slots\leftarrow child\_slots\cap parent\_free\_slots$ $slots\leftarrow\textbf{Select}\ \text{nearest}\ nb\ \text{slots}\ \textbf{from}\ free\_slots$ $child\_slots\leftarrow child\_slots-slots$ Append $slots$ to DAO packet Append $child\_slots$ to DAO packet Transmit DAO packet
10else
       Pass // Retain current situation since there is no slot availability
11
@@
III-B Minimizing 6P Transaction Packets
In the sequenced control packet transmission process of 6TiSCH networks, depicted in Fig. 3, when a node decides to change its parent, it communicates this transition by transmitting a DAO packet to the DAG root and triggers the cell reservation process with the new parent. This entails dispatching a cell reservation request using a 6P transaction as shown in Fig 5. Rather than dedicating an entire packet for this purpose, we propose to judiciously embed the requisite information for cell reservation within the RPL control packet DAO. Since the next hop for each child packet is inherently its parent, all packets, including the modified DAO, will go through the designated parent and hence can extract the embedded information about cell reservation and define its relevant cells. This approach is detailed as follows:
@@ III-B Minimizing 6P Transaction Packets
    Autonomous channel selection: Transmission channel is autonomously determined based on the MAC address of the receiver, thereby enabling nodes to distribute only the slot offset to represent the cell. To avoid interference and benefit from multi-channeling, we adopt the same approach as MSF. Specifically, the current transmission/reception channel ($channel\_{\rm ID}$) is determined based on the Absolute Slot Number (ASN) and the channel offset ($cell_{\rm channel}$) of the current cell as follows:
    $channel\_{\rm ID}=\left(ASN+cell_{\rm channel}\right)\ \%\ nb\_{\rm channels},$ (1)
@@
      Efficient Slot ID encoding: To further enhance the packet size, we introduce an encoding scheme for slot IDs. Typically, 6TiSCH systems limit the slotframe length to 101 slots. However, we propose here to extend the slotframe length to 255 slots, such that the slot ID never exceeds this value. By doing so, we can encode each slot ID into a single byte (8 bits), thus minimizing the slot ID data size injected into the packet.
@@
      Compact free slots notation: Within our proposed approach, the DAO and DIO packets have to convey the node’s free slots. To optimize memory usage, we propose that the node transmits either an ordered list of occupied slots or of free slots, depending on which is shorter and anticipating it with a flag (1 bit). Within the 6TiSCH context, slot “0” is reserved by nodes upon joining the network, hence it will never be included in the ordered list of free slots. This reserved status allows to efficiently utilize slot “0” as the flag for the occupied/free slots list. This flexible method guarantees less memory usage.
    The aforementioned techniques align with the fundamental requirements of low-power and lossy networks (LLNs) since they aim to reduce transmission overhead, save energy, and enhance the network’s throughput. Moreover, they simplify demands’ processing, which is critical for devices operating within constrained environments, such as in 6TiSCH networks.
@@
    TSCH ACK-based optimization of slot allocation: For efficient time slot allocation, we propose a novel method that uses TSCH ACK for immediate confirmation of slot reservations upon DAO packet reception. Doing so bypasses the conventionally incurred delays for reservation confirmation. Leveraging TSCH ACK ensures quicker communication of available time slots, thus reducing the locking duration of time slots (shown as $T_{\rm Lock}$ in Fig. 5) and increasing the network’s flexibility for topological changes. Our technique embeds a list of available slots (identified with integers) directly into the TSCH ACK messages. For the sake of data size reduction, we limit the maximum number of time slots allocated at once to 5 resulting in an additional payload of merely 5 bytes in TSCH ACK. Moreover, to ensure efficient time slot reservation, we assume that each node maintains an up-to-date variable containing its ordered available slots list. This variable is dynamically updated following each slot allocation or de-allocation. Upon receiving a packet, the joining node compares the list of available slots embedded within the packet against its variable of available slots. Hence, the decision to reserve slots is based on the most current network status, which improves the responsiveness and adaptability of the 6TiSCH network to real-time demands.
III-C Reducing DODAG Joining Time
To reduce packet collisions and reduce the nodes’ joining time, we propose that the DIO sender sets temporary time slots or cells, ensuring it is actively listening, and it designates one permanent cell to maintain its ongoing listening state. In scenarios where a DIO is concurrently received by multiple nodes, thus causing potential DAO collisions, the proposed mechanism that uses a combination of temporary and permanent cells can diminish the impact of this issue as follows: 1) The DAO sender temporarily allocates cells until it receives the TSCH ACK of the DAO packet. This strategy intentionally sacrifices idle listening to prioritize energy saving, given that the energy consumed during idle states is notably lower than that needed to manage collisions and trigger packet retransmissions [30]. 2) In the initial phase of network formation where nodes actively compete to join the network and parent switching events are frequent, our method sets a high number of proposed time slots in the DIO to facilitate rapid integration and reduced collision events. As the network stabilizes, we strategically reduce the number of proposed slots to preserve energy.
III-D Scheduling Information Injection

### v2

@@ 2.1.2 Scheduling Function (SF)
Figure 5: Examples of 6P transactions.
2.2 Related Works
The conventional 6TiSCH mechanisms, while foundational, have been scrutinized for potential limitations and areas of enhancement. This section investigates the existing literature for proposed methodologies, achievements, and gaps, thereby contextualizing the motivation and novelty of our work.
Aiming to enhance fundamental aspects of 6TiSCH, particularly the TSCH configuration, the authors of [12] adopted a heterogeneous time slot length to minimize communication latency. However, the use of multiple minimal cells for multi-PHY discovery, dynamic slot duration adaptation, and the added control overhead from frequent slot reconfiguration increase network overhead, memory usage, scheduling complexity, and energy consumption. In [13], the authors introduced a heuristic method for parent selection in multi-PHY TSCH, showing that slot bonding can adapt to different PHY data rates and achieve high Packet Delivery Ratios (PDRs). Regarding 6TiSCH network formation, the study in [14] evaluated the network formation procedure and identified several limitations. Then, the authors proffered several recommendations and established guidelines for parameter adjustments, tailoring parameters to several network topologies. Also, the authors of [15] addressed performance degradation during 6TiSCH network formation. They highlighted the challenges faced by new nodes joining the network, primarily due to the policy of allocating only one shared cell per slotframe for control packet transmission. They particularly emphasized the limitations of prioritizing EBs during network formation. Then, they proposed an improved approach that effectively integrates control packets into the network formation process. In [16], the authors aimed to accelerate the network formation process through the use of active scanning procedures and a Trickled beacon advertising strategy. In their approach, nodes attempting to join transmit beacon request packets, thereby increasing the likelihood of detection by already associated nodes and reducing the overall association time. Although this solution significantly improved network formation, it overlooked energy efficiency as the joining nodes remain in continuous listening mode during the formation phase. In [17], the authors proposed the Enhanced Network Bootstrapping technique using Multi-Channel (EBMC) to improve the network formation. Since EBMC relies on a static channel selection algorithm based on the nodes’ EUI-64 addresses, it suffers from persistent synchronization failures when constant interference affects the assigned channel, thus preventing nodes from joining the network. Furthermore, since topology formation in EBMC does not incorporate dynamic parent selection, inefficient multi-hop routes may occur, resulting in higher end-to-end latency and routing overhead. Despite that multi-channel allocation reduced collision probability, EBMC’s static control packet assignments led to underutilized channel resources during periods of low control packet activity.
Researchers also explored optimizing scheduling functions to enhance network efficiency. Typically, research works that enhance SF fall into three categories: 1) Centralized approaches [18, 19, 20], 2) distributed approaches [21], and 3) autonomous approaches [22]. In [18], the authors proposed a fully centralized schedule for data gathering, where each node has to send the allocation request to the DAG root and receive the answer, thus wasting a considerable amount of energy and time. Alternatively, in [19], a central entity is responsible for managing the scheduling of transmissions in the network and uses a recurrent neural network (RNN)-based model to predict the next transmission slot for each node, aiming to reduce the latency of end-to-end communications. However, a higher control packet overhead is generated. Thus, it imposes a huge burden on the nearest node to the root, and the network formation process takes a long time and consumes a considerable energy, particularly in dense networks. Alternatively, the authors of [20] introduced an Adaptive Static Scheduling (ASS) technique designed to enhance the network’s energy efficiency. In [21], a distributed mechanism to improve routing and scheduling was introduced. However, the allocation of dedicated control cells to minimize collisions led to inefficient resource utilization, particularly in dense and highly mobile deployments. Moreover, the reliance on fixed downlink routes, while beneficial for latency reduction, increased complexity in dynamic environments due to frequent updates and higher control overhead. Finally, the frequent exchange of control messages, route updates, and dedicated cell allocations degraded energy efficiency. In contrast, [22] proposed TESLA, a traffic-aware elastic slotframe adjustment scheme that allows each node in the network to dynamically adjust its slotframe size based on real-time traffic conditions.
Other works focused on optimizing the RPL’s objective function. For instance, authors of [23] introduced a novel parent selection function that accounts for the number of neighbors and traffic load. It optimized the PDR to minimize frequent parent switching. In [24], a Q-learning-based algorithm is proposed to refine the RPL’s objective function, while [25] proposed an early parent switching scheme based on the remaining queue capacity of the parent node. However, its reliance on frequent parent switching and queue monitoring introduces control overhead. Also, it does not account for the fact that a parent’s queue congestion may stem from inefficiencies in the SF rather than true traffic imbalance, potentially leading to suboptimal routing decisions. In addition, [24] suffers from high computational complexity as it uses Q-Learning, which requires frequent Q-value updates, large table storage, and intensive processing, making it less suitable for resource-constrained 6TiSCH networks.
To address limitations inherent to single-layer solutions, cross-layer approaches have been proposed to jointly optimize multi-layer parameters. For instance, authors in [26] proposed a cross-layer scheduling and routing approach that enhances reliability in 6TiSCH networks by integrating alternative parent selection, MPLS-like source routing, and Bounded Delay Packet Control (BDPC). By leveraging multiple alternative parents and implementing a deadline-aware packet forwarding mechanism, their method significantly improved the packet delivery ratio and ensured bounded latency. However, this comes at the cost of higher computational complexity as each node has to process per-packet label switching, maintain multiple parent paths, and track delivery deadlines dynamically. These additional demands increase overhead and are not well-suited for resource-constrained IIoT devices. Also, the approach proposed in [27] improved packet delivery and latency but introduced high computational complexity, increased communication overhead for Q-table sharing, and required careful parameter tuning. In [28], the authors proposed the Improved Minimal Scheduling Function (IMSF) to enhance MSF by dynamically adjusting the number of allocated cells based on queue occupancy and link quality (ETX). This technique improved adaptability to traffic variations and reduced packet loss without generating additional control packets beyond standard 6P transactions. However, IMSF’s dependence on ETX metrics and historical queue data may limit its responsiveness in dynamic network conditions.
As discussed above, existing solutions either introduce additional control overhead, rely on complex computations, or do not adapt to dynamic network conditions, often leading to high delays and energy consumption. Moreover, most works enhanced a single performance metric at the cost of others. Consequently, in this paper, we propose a novel cross-layer approach that accelerates network formation, optimizes scheduling, and improves adaptability while minimizing control overhead, latency, and energy consumption, making it a lightweight yet efficient solution for energy-constrained IIoT environments.
3 Problem Identification
Several operational challenges impede the effectiveness of deploying the 6TiSCH protocol within IIoT networks. They include:
Figure 6: Control packets triggered upon parent switching and joining DODAG [29].
  1. 1.
    Complex Time Slot Allocation for External Nodes: The 6TiSCH protocol experiences difficulties to timely allocate sufficient time slots for nodes not directly part of the DODAG. This issue is critical in scenarios where nodes keep moving, and continuously changing the DODAG’s topology, thus requiring quick slot allocation to maintain connectivity. As depicted in Fig. 6, newly joined nodes, such as $D$, send DAO packets to the DAG root, initiate 6P requests to their parents, and await corresponding 6P responses, before effectively starting to transmit data. However, if a joining node repeatedly fails to send DAO or 6P messages due to unsuccessful transmission or lack of response from the parent, it will eventually lose synchronization and revert to a desynchronized state (See Fig. 4). A lack of timely integration into the network delays operational synchronization and increases energy consumption as new nodes repeatedly attempt to join the network.
  2. 2.
    Excessive Use of Control Packets (6P): The current implementations of the 6TiSCH scheduling function (SF) and routing protocol (RPL) rely heavily on control packets, thus leading to a congested network, which is particularly problematic in large-scale systems A large control packets traffic increases competition for transmission opportunities, which can cause queuing delays, higher latency, and degraded system reliability. This problem is critical in dense networks where simultaneous control packet exchanges occur frequently during network formation, parent switching, and node joining. Fig. 6 illustrates the issue where nodes $L$ and $D$ experience frequent packets exchange during parent switching and DODAG joining, respectively. Both events trigger DAO packets that traverse multiple intermediate nodes before reaching the DAG root, which responds with DAO-ACK messages. Concurrently, the $L$ and $D$ initiate 6P requests to negotiate cell reservations with their new parents, who reply with 6P responses. Such simultaneous DAO and 6P exchanges overload intermediate nodes, create bottlenecks near the DAG root, and delay data delivery.
    (a) $k=5$
    (b) $k=10$
    Figure 7: Impact of free slot availability on first-attempt reservation success probability.
  3. 3.
    Inefficient Management of Time Slot Scheduling: In 6TiSCH networks, when a node initiates a cell reservation process using the 6P protocol, one of the two nodes (either the requester or responder, depending on the transaction type) proposes $k$ candidate cells for potential allocation. To prevent scheduling conflicts, the proposing node temporarily locks these cells until it receives a response. As illustrated in Fig. 5, this cell-locking phase ensures negotiation consistency but may lead to the underutilization of time slots during the reservation process. The locking period can be prolonged in situations involving packet retransmissions or communication delays, further impacting the slotframe efficiency. Moreover, even when the joining node’s slotframe is fully available, the parent node’s slotframe may be heavily occupied due to ongoing communications with other nodes. This issue intensifies during a parent switching where both nodes may have pre-reserved slots, thus restricting the chance of finding common free cells. As slotframe utilization increases, the probability of successfully allocating at least one common free cell on the first attempt decreases, causing multiple reservation attempts and increased setup delays. The probability of successfully reserving at least one common free cell in the first attempt can be given by:
    $P_{success}=1-\left(1-\frac{F_{A}}{C}*\frac{F_{B}}{C}\right)^{k},$ (1)
    where $F_{A}$ and $F_{B}$ are the numbers of free cells at the proposing and responding nodes, respectively, $C$ is the total number of available time slots in the slotframe, and $k$ is the number of proposed cells. Eq.(1) models the probability of at least one successful overlap among the $k$ proposed cells, considering that each time slot has a probability $\left(\frac{F_{A}}{C}*\frac{F_{B}}{C}\right)$ of being free at both nodes. For clarity, we present in Fig. 7 how the success probability increases with the number of free cells on the requester and responder sides. Achieving a success probability above 90% typically requires both nodes to have over 50% of their slotframes free. To evaluate the impact of the number of proposed cells, we analyze scenarios with $k=5$ and $k=10$, with the choice of $k=5$ being consistent with typical configurations described in RFC 9033 [5]. When $k=5$, reaching high $P_{success}$ requires large numbers of free cells on both sides. When $k=10$, achieving a similar $P_{success}$ as for $k=5$ requires a lower number of free cells by 35–40%. Nevertheless, such an improvement comes at the expense of locking more cells during negotiation, potentially causing slotframe underutilization.
  4. 4.
    Inefficient Energy and Time Consumption during Parent Switching: Parent switching in 6TiSCH networks is a critical process that can significantly impact energy consumption and network stability if not properly coordinated with slot allocation. During parent switching, the RPL standard process ignores the new parent’s schedule status, thus potentially leading to an inefficient use of energy and time. This inefficiency stems from the lack of synchronization between the parent selection process handled by RPL and the slot allocation process managed by SF. RPL selects a new parent based solely on routing metrics without verifying if sufficient transmission slots are available. As a result, a node may switch to a parent but fail to send data, causing repeated switching attempts, increased control packet exchanges, and unnecessary energy consumption. Also, queued packets experience increased latency as they wait until the next successful parent switching attempt. This may be due to parent slotframe saturation, simultaneous parent switching by multiple nodes, and delayed slot allocation during cell negotiation. However, in IIoT systems, where devices operate on limited power, the latter’s efficient use is crucial to ensure the system’s longevity and reliability.
  5. 5.
    TSCH Queue Packet Drops due to Cell Shortages: Poor 6TiSCH queue management can significantly degrade network performance, particularly if traffic dynamics are overlooked. Specifically, the TSCH queue can quickly reach capacity, requiring additional transmission cells to accommodate enqueued packets. A mismatch between queue occupancy and the allocated number of cells can lead to packet drops if additional resources are not proactively secured.
In this work, we attempt to address these challenges through the integration of targeted modifications to the standard 6TiSCH protocol, as described in the next section.
4 Proposed Solution
@@
Our work extensively focuses on the intricate interplay between RPL and SF and between SF and packet queue management. The core philosophy of our solution is as follows: 1) Reduce the transmission frequency of control packets upon parent switching; 2) Minimize the number of failed parent switching events that result from cell unavailability; 3) Prevent packet dropping events due to queue saturation; and 4) Improve the jitter and latency performances by slot selection. Our approach targets saving valuable energy resources and mitigating unnecessary medium contention, thus enhancing spatial reuse.
For understanding purposes of our solution, we highlight the following technical terms usage:
@@
  • “DAO-ACK” and “ACK” are two distinct message. DAO-ACK denotes a packet transmitted by the DAG root to acknowledge DAO messages for reliable route registration/de-registration. In contrast, a packet “ACK”, known also as a link-layer or TSCH acknowledgment (shown in Fig. 8), is issued upon packet reception [30].
@@
4.1 Regulated Triggering of Parent Switching
In 6TiSCH networks, parent switching might be challenging. Indeed, the decision to switch to a potentially superior parent is complex, given the node’s lack of awareness regarding the new parent’s cell availability. This uncertainty can lead to inefficient energy usage since the node may continue to control packet transmissions without realizing effective cell allocation. Also, the RPL specifications [31, 6] did not detail the protocol’s behaviour regarding transmission interruption to the former parent during this process. Consequently, increased latency and potential packet loss may occur. To tackle this issue, our solution proposes to modify the DIO packet to carry a list of available slots of the sender. This approach ensures that a node initiates parent switching only when a shared available slot exists, hence improving the efficiency of the process. Algo. 1 below delineates this mechanism, particularly focusing on the node’s decision-making process at the juncture of parent switching.
@@ 4.1 Regulated Triggering of Parent Switching
1
// ------ Initialization ------
2 $nb\_parentSlot\leftarrow\text{Number of slots allocated to parent}$
@@
4
// ------ Parent Selection and Slot Reservation ------
[…]
6
7if $child\_slots\geq min\_nb$ then
@@ 9             $nb\leftarrow min\_nb$
12             $nb\leftarrow nb\_parentSlot$
13             if $nb>max\_nb$ then
@@ 14                   $nb\leftarrow max\_nb$
15
16
17      $free\_slots\leftarrow child\_slots\cap parent\_free\_slots$
18       $slots\leftarrow\text{Select nearest}\ nb\ \text{slots from}\ free\_slots$
19       $child\_slots\leftarrow child\_slots-slots$
20       Append $slots$ to DAO packet
21       Append $child\_slots$ to DAO packet
22       Transmit DAO packet
23
24else
25       Pass
       // Retain current situation since there is no slot availability
26
27
@@ 25       Pass
4.2 Minimizing 6P Transaction Packets
In the sequenced control packet transmission process of 6TiSCH networks, depicted in Fig. 3, when a node decides to change its parent, it communicates this transition by transmitting a DAO packet to the DAG root and triggers the cell reservation process with the new parent. This entails dispatching a cell reservation request using a 6P transaction as shown in Fig.5.
Rather than dedicating the entire packet and the time slot for this purpose, we propose to judiciously embed the requisite information for cell reservation within the RPL control packet DAO. Since the next hop for each child packet is inherently its parent, all packets, including the modified DAO, will go through the designated parent and hence can extract the embedded information about cell reservation and define its relevant cells. This approach is detailed as follows:
@@ 4.2 Minimizing 6P Transaction Packets
    Autonomous channel selection: Transmission channel $cell_{\rm channel}$ is autonomously determined based on the MAC address of the receiver using a transformation function $f$, thereby enabling nodes to distribute only the slot offset to represent the cell. $cell_{\rm channel}$ is defined during the reservation using the equation below, which follows the same method used to determine the channel of an autonomous cell:
    $cell_{\rm channel}=f(MAC_{\rm address})\ \%\ nb\_{\rm channels},$ (2)
@@
    $channel\_{\rm ID}=\left(ASN+cell_{\rm channel}\right)\ \%\ nb\_{\rm channels}.$ (3)
@@
      Slot ID encoding: We propose to implement an 8-bit encoding scheme for slot IDs, thus minimizing the size of the slot ID list injected into the packet.
@@
      Compact free slots notation: Within our proposed approach, the DAO and DIO packets have to convey the node’s free slots. To optimize memory usage, we propose that the node transmits either an ordered list of occupied slots or of free slots, depending on which is shorter, and anticipating it with a flag (1 bit). Within the 6TiSCH context, slot “0” is reserved by nodes upon joining the network, hence, it will never be included in the ordered list of free slots. This reserved status allows us to efficiently utilize slot “0” as the flag for the occupied/free slots list. Since each joined node reserves at least three cells (for Minimal cell, Autonomous Rx, and Autonomous Tx), The largest size of the list that will be injected is
      $L_{Length}=\frac{Slotframe_{Length}}{2}-3,$ (4)
      where $Slotframe_{Length}$ is the size of the slotframe (in number of time slots). This flexible method guarantees less memory usage.
    The aforementioned techniques align with the fundamental requirements of LLNs since they aim to reduce transmission overhead, save energy, and enhance the network’s throughput. Moreover, they simplify demand processing, which is critical for devices operating within constrained environments, such as in 6TiSCH networks.
@@
    TSCH ACK-based optimization of slot allocation: For efficient time slot allocation, we propose a novel method that uses TSCH ACK for immediate confirmation of slot reservations upon DAO packet reception. Doing so bypasses the conventionally incurred delays for reservation confirmation. Leveraging TSCH ACK ensures quicker communication of available time slots, thus reducing the locking duration of time slots (shown as $T_{\rm Lock}$ in Fig. 5) and increasing the network’s flexibility for topological changes. Our technique embeds a list of selected slots (identified with integers) directly into the TSCH ACK messages. We limit the maximum number of time slots allocated at once to $5$ resulting in an additional payload of merely $5$ bytes in TSCH ACK. Moreover, to enable fast and timely slot selection within the strict time constraints of TSCH, we assume that each node maintains an up-to-date variable containing its ordered available slots list. This variable is dynamically updated following each slot allocation or de-allocation. Upon receiving the DAO packet, the node compares the list of available slots embedded within the packet against its variable of available slots. Hence, the decision to reserve slots is based on the most current network status, which improves the responsiveness and adaptability of the 6TiSCH network to real-time demands.
4.3 Reducing DODAG Joining Time
To reduce packet collisions during network formation and the nodes’ joining time, we propose that the DIO sender sets temporary time slots, ensuring it is actively listening, and it designates one permanent cell to maintain its ongoing listening state. In this way, nodes that receive the DIO packets and attempt to join the DODAG, transmit the DAO packets during the proposed slots, with the channel defined according to Eq.(3). When multiple nodes attempt to send the DAO packet to the DIO sender, they compete for multiple available cells instead of a single cell, thus reducing contention and improving transmission efficiency. To streamline network formation, the proposed mechanism dynamically adjusts the allocation of temporary slots. Initially, a high number of temporary slots, defined as “Number of proposed slots per DIO”, are made available for allocation to accelerate node integration and reduce joining time. These slots are updated with each DIO transmission, ensuring efficient resource management. The DIO sender remains in listening mode during these slots for a duration of “DIO cells duration”, after which it listens only to the permanent slot. The receiver of the DIO packet selects a portion of these slots to send the DAO packet, ensuring controlled channel access and minimizing contention.
4.4 Scheduling Packet Update